# Wearable Accelerometer Foundation Models for Health via Knowledge Distillation

### v1

@@ 6 Discussion, limitations and future work
Here, we demonstrated that a single accelerometry encoder can predict heart rate, heart rate variability, demographic variables and a wide array of downstream health targets, i.e., as a foundation model. This was done with a fully unsupervised representational knowledge distillation framework to transfer knowledge from high-fidelity to low-fidelity biosignals. We showed that our knowledge distillation framework is not unique to the teacher/student architecture, or teacher pre-training method, and it results in near-perfect alignment of PPG and accelerometry embeddings as quantified by our retrieval analysis. In addition, it improves the quality of the representations in the low-fidelity biosignal compared to self-supervised and supervised baselines, while maintaining label efficiency for downstream targets. We also showed that cross-modal distillation can simultaneously enable compressing the student model while transferring knowledge across modalities. We also performed ablations with respect to several training choices. Our work primarily focuses on accelerometry signals during low-motion and sedentary periods (Section 4.1), where accelerometer captures ballistocardiogram and therefore minute cardiovascular-related information useful for health targets (Inan et al., 2015; Kim et al., 2016). Future work can investigate models that take slower-scale activity metrics on wearable devices such as steps, speed, sleep, and slow changes in accelerometry, as well as minute changes in accelerometry at sedentary settings to improve the performance for downstream targets (Hallgrímsson et al., 2018; Ni et al., 2019; Spathis et al., 2021). In addition, another interesting area of investigation for future work could be experimenting with modality specific augmentations (Qian et al., 2022; Demirel & Holz, 2023). One caveat of our work is that it currently supports two modalities, while future work can consider statistical objectives for mutual information maximization, when there are more than one teacher or student modalities (Shidani et al., 2024), or by binding all student modalities to a single teacher modality (Girdhar et al., 2023). Another caveat is that while this work can be used for knowledge transfer or retrieval of the high-fidelity modality embeddings, it does not provide a generative model across modalities (Sarkar & Etemad, 2020); future work can consider recent techniques to incorporate generative capabilities using unified encoder and decoder Transformers (Mizrahi et al., 2023; Meta, 2024). Ideally, future work can also consider modeling other modalities such as text, images and videos to leverage information from these other input sources and weakly supervise biosignal representations, similar to prior work for modeling accelerometry during motion and other modalities such as video and text (Moon et al., 2022; Tan et al., 2023).
@@ Acknowledgments
Reproducibility statement
The aggregated data that support the findings of this study can be made available on request from the corresponding author. Request for data will be evaluated and responded to in a manner consistent with the specific language in the study protocol and informed consent form (ICF). Based on the language within the IRB approved ICF for the Apple Heart and Movement Study, we are unable to share sensor data collected in the study. Similarly, code for all data analyses may be available upon request from the corresponding author. Requests for code will be evaluated and responded to in a manner consistent with policies intended to protect participant confidentiality and language in the study protocol and ICF.
@@ A.2 Dataset and evaluations
Linear probing for age, BMI, biological sex, and health targets from AHMS survey questions: We perform linear probing for predicting self-reported age, body mass index (BMI), biological sex (sex assigned at birth), and health targets from survey questions for participants in AHMS. During AHMS, participants fill out multiple questionnaires containing various questions regarding their historical health record and demographics (Truslow et al., 2024; Abbaspourazad et al., 2024). The response to these questions are usually in form of ‘yes’ or ‘no’ for whether the participant has had a history of a health condition (e.g., asthma), or whether they take specific medications (e.g., anti-depressants), or regarding their lifestyle habits (e.g., smoking). For the classification tasks, we use ridge regression to predict scores for binarized targets (0/1) and we quantify the performance with area under curve of receiver’s operating curve (AUC). For biological sex, we classify male versus female, and for health targets we classify ‘yes’ versus ‘no’. For regression tasks (age and BMI), we use ridge regression to predict continuous targets and we use mean absolute error to quantify performance. Age is reported in years and BMI is reported in $\text{kg}/\text{m}^{2}$. In all these subject-related targets that do not vary from segment to segment, we perform the linear probing at participant granularity: we mean-aggregate all the embeddings associated to each participant so that each participant contributes one and only one sample in the downstream training/evaluation. Similar to the heart rate and heart rate variability linear probing, the downstream training/evaluation splits for these targets are stratified based on participants such that the evaluation split’s participants have no overlap with those in the training split of the linear probing or pre-training.
@@ A.3 Results and ablation studies
Ablation on choice of positive pairs in teacher pre-training with contrastive learning: We selected the positive pairs as two augmented views of the same sample, as opposed to two different segments of the same participants proposed in (Abbaspourazad et al., 2024). This choice was made to enforce the encoders to contain more segment-level information necessary for the main downstream targets used in this study (heart rate and heart rate variability). Appendix Table 14 demonstrates the performance of heart rate, SDNN and RMSSD for the “PPG-CL” trained with two different positive pair selection strategies.
Ablation on number of PPG channels in teacher pre-training: We performed an ablation about the effect of the number of PPG channels in teacher pre-training on downstream evaluations in Table 15, where we only kept one of the PPG channels for modeling and observed a drop in performance.
@@
Ablation on augmentations: In addition to comparing knowledge distillation with and without augmentations (Section 5.5), we studied the importance of individual augmentations during knowledge distillation. To this end, we performed the knowledge distillation from PPG to accelerometry, where we only kept one of the augmentation functions during distillation (applied in every forward pass), one at a time. This was done while maintaining all other training choices the same to control for the effect of augmentations. We observed that: 1) our stochastic augmentation module achieved the highest accuracy (Appendix Table 17), likely because it captures more diverse range of distortions during training, 2) among the individual augmentations, ”add Gaussian noise” and ”Cut out” had the highest importance, while ”Time warp” had the least importance.
@@
Ablation on impact of $\boldsymbol{\lambda}$ in the multi-modal contrastive objective $\mathbf{\lambda}$: We studied the impact of $\lambda$ in Equation 2 for our representational knowledge distillation. We observed that while the improvements of accelerometry embeddings via distillation were robust to $\lambda$, higher values of $\lambda$ ($\lambda=\{0.75,1\}$) were the most optimal (Appendix Table 18), indicating that keeping PPG embeddings as anchor embeddings provided the most knowledge transfer, perhaps due to the fact that PPG is the higher-fidelity modality.
Ablation on initialization of the PPG encoder in multi-modal pre-training: In multi-modal pre-training mentioned in Section 5.5, even if we initialize the PPG encoder with PPG-MAE (Table 19), the downstream predictions are not as good as knowledge distillation in Table 3, indicating the importance of freezing the teacher encoder.
@@ A.4 Discussion
Table 19: Ablation on not freezing the PPG encoder in cross-modal knowledge distillation in Table 3.

### v2

@@ 6 Discussion, limitations and future work
Here, we demonstrated that a single accelerometry encoder can predict heart rate, heart rate variability, demographic variables and a wide array of downstream health targets, i.e., is a foundation model. This was done with a fully unsupervised knowledge distillation framework to transfer knowledge from PPG to accelerometry to improve performance. We showed that our observations are not unique to the teacher/student architecture, or teacher pre-training method, and we can achieve near-perfect alignment of PPG and accelerometry embeddings. In addition, we showed improvements in the quality of the representations compared to self-supervised and supervised baselines, while maintaining label efficiency for downstream targets. We also showed that cross-modal knowledge distillation can enable compressing the student model size. Our work primarily focuses on accelerometry signals during low-motion and sedentary periods (Section 4.1), where accelerometer captures ballistocardiogram and therefore minute cardiovascular-related information useful for health targets (Inan et al., 2015; Kim et al., 2016). Future work can investigate models that take slower-scale activity metrics on wearable devices such as steps, speed, sleep, and slow changes in accelerometry, as well as minute changes in accelerometry at sedentary settings to improve the performance for downstream targets (Hallgrímsson et al., 2018; Ni et al., 2019; Spathis et al., 2021; Xu et al., 2024). In addition, another interesting area of investigation for future work could be experimenting with modality specific augmentations (Qian et al., 2022; Demirel & Holz, 2023). One caveat of our work is that it currently supports two modalities, while future work can consider statistical objectives for mutual information maximization, when there are more than one teacher or student modalities (Shidani et al., 2024), or by binding all student modalities to a single teacher modality (Girdhar et al., 2023). Another caveat is that while this work can be used for knowledge transfer or retrieval of the high-fidelity modality embeddings, it does not provide a generative model across modalities (Sarkar & Etemad, 2020); future work can consider recent techniques to incorporate generative capabilities using unified encoder and decoder Transformers (Mizrahi et al., 2023; Meta, 2024). Ideally, future work can also consider modeling other modalities such as text, images and videos to leverage information from these other input sources and weakly supervise biosignal representations, similar to prior work for modeling accelerometry during motion and other modalities such as video and text (Moon et al., 2022; Tan et al., 2023).
Impact Statement
This work advances machine learning in digital health by enabling the extraction of health insights from accelerometer data commonly available in wearable devices. This could increase the accessibility of health monitoring and improve health outcomes through early detection of health conditions. However, we acknowledge that such models may exacerbate equity gaps between individuals with access to wearable devices and those without. It is also important to address potential ethical concerns, such as privacy and potential biases in model outcomes. We encourage continued research into interpretability methods suited to foundation models in healthcare contexts. Responsible development and deployment of these models necessitates careful consideration of these factors to ensure equitable and beneficial outcomes for all individuals.
@@ A.2 Dataset and evaluations
Linear probing for age, BMI, biological sex, and health targets from AHMS survey questions: We perform linear probing for predicting self-reported age, body mass index (BMI), biological sex (sex assigned at birth), and health targets from survey questions for participants in AHMS. During AHMS, participants fill out multiple questionnaires containing various questions regarding their historical health record and demographics (Truslow et al., 2024; Abbaspourazad et al., 2024). The response to these questions are usually in form of ‘yes’ or ‘no’ for whether the participant has had a history of a health condition (e.g., asthma), or whether they take specific medications (e.g., anti-depressants), or regarding their lifestyle habits (e.g., smoking). For the classification tasks, we use ridge regression to predict scores for binarized targets (0/1) and we quantify the performance with area under curve of receiver’s operating curve (AUC). For biological sex, we classify male versus female, and for health targets we classify ‘yes’ versus ‘no’. For regression tasks (age and BMI), we use ridge regression to predict continuous targets and we use mean absolute error to quantify performance. Age is reported in years and BMI is reported in $\text{kg}/\text{m}^{2}$. Given that these targets do not vary from segment to segment, we perform the linear probing at participant granularity: we mean-aggregate all the embeddings associated to each participant so that each participant contributes one and only one sample in the downstream training/evaluation. Similar to the heart rate and heart rate variability linear probing, the downstream training/evaluation splits for these targets are stratified based on participants such that the evaluation split’s participants have no overlap with those in the training split of the linear probing or pre-training.
@@ A.3 Results and ablation studies
Accelerometry model compression via knowledge distillation: We demonstrate that we can perform model compression for the student accelerometry encoder in Figure 4, as explained in Section 5.4.
@@
Ablation on choice of positive pairs in teacher pre-training with contrastive learning: We selected the positive pairs as two augmented views of the same sample. This was done to enforce the encoders to contain more segment-level information necessary for the main downstream targets presented in this study including heart rate and heart rate variability. There are other choices of positive pair selections in prior work, for example participant-level positive pair selection (Abbaspourazad et al., 2024; Pillai et al., 2024). Appendix Table 14 demonstrates the performance of heart rate, SDNN and RMSSD for the “PPG-CL” trained with these two different positive pair selection strategies.
Ablation on number of PPG channels in teacher pre-training: We performed an ablation about the effect of the number of PPG channels in teacher pre-training on downstream evaluations in Table 15, where we only kept one of the PPG channels for modeling and observed a drop in performance. This demonstrates the importance of modeling multi-channel PPG signals.
@@
Ablation on augmentations: In addition to comparing knowledge distillation with and without augmentations, we studied the importance of individual augmentations during knowledge distillation (Section 5.5). To this end, we performed the knowledge distillation from PPG to accelerometry, where we only kept one of the augmentation functions during distillation (applied in every forward pass), one at a time. This was done while maintaining all other training choices the same to control for the effect of augmentations. We observed that: 1) our stochastic augmentation module achieved the highest accuracy (Appendix Table 17), likely because it captures more diverse range of distortions during training, 2) among the individual augmentations, ”add Gaussian noise” and ”Cut out” had the highest importance, while ”Time warp” had the least importance.
@@
Ablation on impact of $\boldsymbol{\lambda}$ in the cross-modal knowledge distillation $\mathbf{\lambda}$: We studied the impact of $\lambda$ in Equation 1 for our representational knowledge distillation. We observed that while the improvements of accelerometry embeddings via distillation were robust to $\lambda$, higher values of $\lambda$ ($\lambda=\{0.75,1\}$) were the most optimal (Appendix Table 18), indicating that keeping PPG embeddings as anchor embeddings provided the most knowledge transfer, perhaps due to the fact that PPG is the higher-fidelity modality.
Ablation on initialization of the PPG encoder and $\boldsymbol{\lambda}$ in comparison to simultaneous multi-modal pre-training: In multi-modal pre-training mentioned in Section 5.5, even if we initialize the PPG encoder with PPG-MAE (row 2 of Table 19), the downstream predictions are not as good as knowledge distillation (row 3 of Table 3). We also observed that the conclusions remained in tact with different values of $\lambda$ in the contrastive learning objective in Equation 1.
@@ A.4 Discussion
Embedding Top-1 Acc. $\uparrow$ Mean Rank $\downarrow$
Accel-KD via PPG-MAE 99.17 (0.23) 1.02 (0.01)
Accel-MAE + Procrustes align. 0.31 (0.04) 2311.50 (60.74)
Chance-level performance 0.01 9551.64
@@
Name Accel-KD via PPG-MAE Accel-KD via PPG-CL Accel-MAE Accel-CL
ACE-inhibitors 0.802 (***) 0.794 0.731 0.791
Active alcohol user 0.681 (***) 0.675 0.616 0.665
Active smoker 0.810 (***) 0.801 0.735 0.784
Afib 0.816 (***) 0.804 0.765 0.798
Allergy 0.652 (***) 0.648 0.619 0.644
Anti-anxiety 0.713 (***) 0.707 0.641 0.696
Anti-psychotics 0.796 (***) 0.785 0.705 0.767
Anticoagulants 0.818 (***) 0.809 0.759 0.801
Antidepressants 0.795 (***) 0.782 0.685 0.761
Antiplatelets 0.784 (***) 0.781 0.732 0.776
Anxiety 0.767 (***) 0.759 0.679 0.747
Artery disease 0.880 (*) 0.869 0.822 0.873
Arthritis 0.781 (***) 0.773 0.733 0.774
Asthma 0.634 (***) 0.630 0.596 0.621
Beta-blockers 0.759 (***) 0.747 0.690 0.736
Blood pressure 0.798 (***) 0.789 0.732 0.787
Blood pressure med. 0.710 (***) 0.697 0.651 0.694
Calcium-channel blockers 0.772 (***) 0.759 0.703 0.757
Cancer 0.800 (***) 0.791 0.743 0.793
Chemotherapy 0.735 (***) 0.704 0.626 0.714
Cholesterol 0.755 (***) 0.747 0.703 0.746
Chronic bronchitis 0.725 (***) 0.725 0.683 0.714
Depression 0.740 (***) 0.735 0.665 0.722
Diabetes 0.829 (***) 0.818 0.767 0.810
Diuretics 0.756 (***) 0.747 0.701 0.743
Hearing 0.719 (***) 0.713 0.676 0.709
Heart attack 0.835 (*) 0.832 0.771 0.831
Heart disease 0.857 (***) 0.845 0.801 0.843
Heart failure 0.857 (n.s.) 0.838 0.789 0.855
Heart rhythm 0.678 (***) 0.664 0.634 0.663
Hip/Knee 0.844 (*) 0.842 0.790 0.841
Kidney 0.694 (***) 0.687 0.646 0.678
Liver 0.729 (***) 0.713 0.615 0.696
Lower back 0.685 (***) 0.681 0.651 0.674
Neck disorder 0.724 (***) 0.717 0.676 0.714
Neuropathy 0.802 (***) 0.793 0.747 0.791
Opioid painkillers 0.769 (***) 0.763 0.667 0.748
Osteoporosis 0.854 (***) 0.849 0.807 0.851
Pacemaker 0.910 (***) 0.884 0.835 0.885
Painkillers 0.602 (***) 0.599 0.578 0.597
Sleep apnea 0.798 (***) 0.798 0.729 0.783
Sleep medication 0.673 (***) 0.661 0.622 0.652
Stroke or TIA 0.790 (n.s.) 0.779 0.743 0.789
Thyroid 0.750 (***) 0.746 0.712 0.743
Urinary 0.799 (***) 0.790 0.748 0.787
Vision 0.657 (***) 0.655 0.627 0.651
@@
Heart rate (BPM) $\downarrow$ SDNN (ms) $\downarrow$ RMSSD (ms) $\downarrow$
Accel-CL (EfficientNet) 2.33 (2.54) 9.93 (10.54) 10.95 (12.39)
PPG-CL (EfficientNet) 1.20 (1.28) 7.17 (7.55) 8.02 (8.71)
Accel-KD via PPG-CL (EfficientNet) 1.37 (1.58) 7.16 (8.04) 8.42 (9.57)
Accel-CL (Transformer) 2.28 (2.48) 9.81 (10.43) 10.87 (12.09)
PPG-CL (Transformer) 1.13 (1.20) 7.23 (7.58) 8.08 (8.73)
Accel-KD via PPG-CL (Transformer) 1.41 (1.59) 7.17 (7.92) 8.67 (9.67)
@@
Eval. $\downarrow$ Participant-level positive pairs Segment-level positive pairs
Heart rate (BPM) 2.51 1.21
SDNN (ms) 12.24 6.58
RMSSD (ms) 11.76 8.40
@@
Eval. $\downarrow$ PPG-MAE w/ 1 PPG channels PPG-MAE w/ 4 PPG channels
Heart rate (BPM) 0.39 0.34
SDNN (ms) 5.04 4.46
RMSSD (ms) 7.30 6.34
@@
Eval. $\downarrow$ PPG-MAE (12.7M) PPG-MAE (6.3M)
Heart rate (BPM) 0.34 0.34
SDNN (ms) 5.05 4.46
RMSSD (ms) 7.52 6.34
@@
Eval. $\downarrow$ W/o augs. W augs. Cut out Gauss. noise Mag. warp Channel perm. T. warp
Heart rate (BPM) 1.76 1.21 1.38 1.29 1.76 1.79 1.84
SDNN (ms) 6.98 6.58 6.96 6.66 6.97 6.77 10.47
RMSSD (ms) 8.82 8.40 8.51 8.84 8.96 8.67 12.83
@@
Eval. $\downarrow$ $\boldsymbol{\lambda=0}$ $\boldsymbol{\lambda=0.25}$ $\boldsymbol{\lambda=0.5}$ $\boldsymbol{\lambda=0.75}$ $\boldsymbol{\lambda=1}$
Heart rate (BPM) 1.25 1.26 1.25 1.19 1.21
SDNN (ms) 6.80 6.69 6.77 6.80 6.58
RMSSD (ms) 8.78 8.59 8.68 8.78 8.40
Table 19: Ablation on $\lambda$ and not freezing the PPG encoder in simultaneous multi-modal contrastive learning in Table 3.
Pre-training framework Eval. $\downarrow$ $\boldsymbol{\lambda=0}$ $\boldsymbol{\lambda=0.5}$ $\boldsymbol{\lambda=1}$
Simultaneous multi-modal contrastive learning (PPG encoder initialized randomly) Heart rate (BPM) 2.81 2.59 2.36
SDNN (ms) 11.01 10.37 9.68
RMSSD (ms) 12.55 11.90 11.30
Simultaneous multi-modal contrastive learning (PPG encoder initialized with “PPG-MAE” weights) Heart rate (BPM) 2.73 2.61 2.23
SDNN (ms) 10.97 10.29 9.63
RMSSD (ms) 12.56 11.89 11.06
Cross-modal knowledge distillation (ours, Accel-KD via PPG-MAE) Heart rate (BPM) 1.25 1.25 1.21
SDNN (ms) 6.80 6.77 6.58
RMSSD (ms) 8.78 8.68 8.40
@@
Target label Medical condition N (test)
Heart attack Heart attack (myocardial infarction) 26,806
Heart disease Coronary heart disease or angina pectoris 26,584
Blood pressure High blood pressure (hypertension) 26,326
Stroke or TIA Stroke (cerebral hemorrhage, cerebral thrombosis) or transient ischemic attack (ministroke) 26,805
Afib Atrial fibrillation 26,342
Heart rhythm Heart rhythm problem other than atrial fibrillation 26,068
Pacemaker Pacemaker 26,916
Artery disease Peripheral artery disease 26,429
Heart failure Heart failure 26,843
Diabetes Diabetes 26,661
Cholesterol High cholesterol 26,238
Arthritis Arthritis 26,429
Hip/Knee Hip or knee replacement 26,948
Lower back Low back disorder or other chronic back defect 26,480
Neck disorder Neck disorder or other chronic neck defect 26,628
Osteoporosis Osteoporosis 26,554
Asthma Asthma 26,762
Chronic bronchitis Chronic bronchitis, chronic obstructive pulmonary disease, or emphysema 26,707
Allergy Rhinitis, hay fever, eye inflammation, dermatitis, food allergy or other allergy (allergic asthma excluded) 26,626
Kidney Kidney problems 26,640
Thyroid Thyroid disease 26,502
Cancer Cancer 26,783
Liver Cirrhosis of the liver 26,784
Urinary Urinary incontinence 26,728
Neuropathy Neuropathy 26,370
Depression Depression 26,110
Anxiety Anxiety disorder 25,989
Hearing Do you have hearing loss? 25,633
Vision Do you have vision loss? 25,895
@@
Target label Medications N (test)
ACE-inhibitors ACE-inhibitors or ARBs (for blood pressure) such as captopril, enalapril, lisinopril, losartan, ramipril, or valsartan 11,043
Anti-anxiety Anti-anxiety aids such as alprazolam (Xanax®), clonazepam (Klonopin®), clorazepate (Tranxene®), diazepam (Valium®), or lorazepam (Ativan®) 15,876
Anti-psychotics Anti-psychotics such as haloperidol (Haldol®), aripiprazole (Abilify®), risperidone (Risperdal®), quetiapine (Seroquel®), olanzapine (Zyprexa®), clozapine (Clozaril®), or lurasidone (Latuda®) 15,914
Anticoagulants Anticoagulants (blood thinners) such as warfarin (Coumadin®), apixaban (Eliquis®), betrixaban (Bevyxxa®), dabigatran (Pradaxa®), edoxaban (Lixiana®), or rivaroxaban (Xarelto®) 15,906
Antidepressants Antidepressants such as amitriptyline (Elavil®), bupropion (Wellbutrin®), citalopram (Celexa®), duloxetine (Cymbalta®), escitalopram (Lexapro®), fluoxetine (Prozac®), paroxetine (Paxil®), mirtazapine (Remeron®), sertraline (Zoloft®), or venlafaxine (Effexor®) 15,919
Antiplatelets Antiplatelets (blood thinners) such as aspirin, clopidogrel (Plavix®), prasugrel (Effient®), or ticagrelor (Brilinta®) 15,891
Beta-blockers Beta-blockers (for blood pressure or heart rhythm) such as atenolol (Tenormin®), bisoprolol (Zebeta®), carvedilol (Coreg®), labetalol, metoprolol (Lopressor®, Toprol-XL®), nadolol (Corgard®), nebivolol (Bystolic®), propranolol (Inderal®), or sotalol (Betapace®) 15,868
Blood pressure med. Other medications for lowering blood pressure such as clonidine, hydralazine, minoxidil, or sacubitril/valsartan (Entresto®) 15,835
Calcium-channel blockers Calcium-channel blockers (for blood pressure or heart rhythm) such as amlodipine (Norvasc®), diltiazem, or verapamil 15,812
Chemotherapy Certain types of chemotherapy such as carboplatin, cisplatin, oxaliplatin, vincristine, or vinblastine 11,084
Diuretics Diuretics (water pills) such as chlorthalidone, furosemide (Lasix®), hydrochlorothiazide, or spironolactone 15,929
Opioid painkillers Opioid painkillers such as codeine, fentanyl, hydrocodone, hydromorphone (Dilaudid®), meperidine (Demerol®), morphine, oxycodone, Percocet®, or Vicodin® 15,941
Painkillers Non-steroidal anti-inflammatories (painkillers) such as aspirin, celecoxib (Celebrex®), diclofenac (Cambia®), ibuprofen (Motrin®/Advil®), or naproxen (Aleve®) 15,919
Sleep medication. Sleeping aids such as eszopiclone (Lunesta®), zaleplon (Sonata®), or zolpidem (Ambien®) 15,892
@@
# Question Answer choices
Q1 In your entire life, have you had at least 1 drink of any kind of alcohol, not counting small tastes or sips? ’Yes’/’No’/’I don’t know’/’I prefer not to answer’
Q1b [If yes to Q1] How often did you have a drink containing alcohol in the past year? ’Never’/’Monthly or less’/’Two to four time a month’/’Two to three times a week’/’Four or more times a week’/’I prefer not to answer’
Q1c [If yes to Q1] On a typical day when you drink, how many drinks do you have? ’1 or 2’/’3 or 4’/’5 or 6’/’7 to 9’/’10 or more’/’I prefer not to answer’
Q2 Have you smoked at least 100 cigarettes in your entire life? ’Yes’/’No’/’I don’t know’/’I prefer not to answer’
Q2b [If Yes or Do not know to Q2] Do you now smoke cigarettes every day, some days, or not at all? ’Every day’/’Some days’/’Not at all’/’I prefer not to answer’
@@
Target label Logic N (test)
Active alcohol user ’Yes’: answer to Q1 is ’Yes’ and answer to Q1b is ’Two to three times a week’/’Four or more times a week’ 23,472
’No’: answer to Q1 is ’No’, or answer to Q1b is ’Never’/’Monthly or less’/’Two to four time a month’
Active smoker ’Yes’: answer to Q2 is ’Yes’ and answer to Q2b is ’Every day’/’Some days’ 23,452
’No’: answer to Q2 is ’No’, and answer to Q2b is ’Not at all’